# Comparing Hyper-optimized Machine Learning Models for Predicting Efficiency Degradation in Organic Solar Cells

### v2

@@ Abstract
This work presents a set of optimal machine learning (ML) models to represent the temporal degradation suffered by the power conversion efficiency (PCE) of polymeric organic solar cells (OSCs) with a multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/
Al. To that aim, we generated a database with 996 entries, which includes up to 7 variables regarding both the manufacturing process and environmental conditions for more than 180 days. Then, we relied on a software framework that brings together a conglomeration of automated ML protocols that execute sequentially against our database by simply command-line interface. This easily permits hyper-optimizing and randomizing seeds of the ML models through exhaustive benchmarking so that optimal models are obtained. The accuracy achieved reaches values of the coefficient determination (R2) widely exceeding 0.90, whereas the root mean squared error (RMSE), sum of squared error (SSE), and mean absolute error (MAE) $\lesssim$1of the target value, the PCE. Additionally, we contribute with validated models able to screen the behavior of OSCs never seen in the database. In that case, R2$\sim$0.96-0.97 and RMSEsimilar-$\sim$1%, thus confirming the reliability of the proposal to predict. For comparative purposes, classical Bayesian regression fitting based on non-linear mean squares (LMS) are also presented, which only perform sufficiently for univariate cases of single OSCs. Hence they fail to outperform the breadth of the capabilities shown by the ML models. Finally, thanks to the standardized results offered by the ML framework, we study the dependencies between the variables of the dataset and their implications for the optimal performance and stability of the OSCs. Reproducibility is ensured by a standardized report altogether with the dataset, which are publicly available at Github [^1] https://github.com/davidumh11.
@@
Organic solar cell, Degradation, P3HT:PCBM, PEDOT:PSS, Machine Learning, Regression.
journal: Engineering Applications of Artificial Intelligence
@@
[label3]organization=Institute of Chemical Synthesis and Homogeneous Catalysis (ISQCH), CSIC, University of Zaragoza,addressline=Pedro Cerbuna 12, city=Zaragoza, postcode=50009, country=Spain
@@ 1 Introduction
Nowadays electricity consumption has evolved exponentially over the last three decades, increasing from a total estimate of 15107.8 TWh to 28661 TWh between 2000 and 2022, respectively Rit2023 . One of the immediate consequences has been reflected in the price of this energy. Simultaneously, in the technological realm, there has been a notable shift in electricity generation towards renewable sources. Particularly, the maturity of conventional photovoltaic (PV) cells and the emergence of new organic and hybrid materials have made possible to get significant improvements in PV generation and its power conversion efficiency (PCE) NREL . However, despite the fact that the global consumption of PV energy increases from 1.08 TWh to 1310.02 TWh between 2000 and 2022 Rit2023 , its generation still falls short of meeting current demand. Hence, there are many challenges and open lines for improving the generation and efficiency of PV cells.
@@
In this context, organic solar cells (OSCs) have emerged as a promising alternative to silicon-based solar cells since, amongst other benefits, their production involves low-temperature manufacturing methods Seo2015 , reducing their carbon footprint. Their appeal also lies in the ease of processing, low cost, flexibility, and lightweight Yeh2013 . Moreover, the evolution of their PCEs has been much superior to classical technologies and it has been on par with other considered emerging technologies, presenting efficiencies that exceed values of 20% veintePCE . However, the commercialization of OSCs still faces various challenges, including stability and efficiency. Nevertheless, certain cost studies PVCost1 ; PVCost2 indicate that once these obstacles are overcome, OSCs can be manufactured at a cost lower than 1 dollar per peak Watt Bra2005 . This fact suggests that the optimization of the manufacturing procedure of these organic devices would certainly lead to reinforce their benefits, but also, and most importantly, their sustainability and stability to operate long-term feasibly. All in all, the potential capabilities of artificial intelligence (AI) protocols, through particular machine learning (ML)-based models, envisages great achievements in this sense.
@@
The integration of ML before transferring to industrial implementation it is being extensively tested these days. For instance, within the field of cheminformatics, it is of paramount importance in the design of new materials. Numerous studies have demonstrated the potential of deep learning algorithms to analyze complex chemical data, thereby accelerating the process of discovering, for example, drugs, as well as identifying promising compounds with improved properties Lus2013 . Some of the early studies Han2013 ; Rup2012 laid the foundations for the application of ML models in predicting molecular properties, highlighting their great strengths. More recently, the deployment of AI techniques with high-performance prediction methods has enabled the emergence of software platforms capable of examining extensive chemical databases to identify potential materials with significant properties Unt2014 ; Wall2015 . A relevant database, which moved forward from general chemical structures to the inclusion of PV features was the project presented in Joh2011 . This work allowed to obtain custom predictive models that characterize the electrical behavior of PV devices in terms of parameters such as voltage, current, conversion efficiency, short-circuit current, open-circuit voltage, etc. Eib2021 ; Ser2016 ; Tol2019 ; Lop2016 . Thus there is an endless list of mathematical models mostly supported by ML and deep learning-based engines that combine different techniques, such as regression and statistical inference Moo2022 ; Eib2021 ; Mal2021 ; Seo2019 ; Sun2019 to optimize different target features of optoelectronic devices Miy2021 ; Mah2022a ; Mah2022b . Similarly, the growth of proposals with eco-design characteristics and AI support for the production of sustainable and durable energy generating materials is positively evolving Vel2023 ; Ju2018 ; Mah2023 . More specifically, some studies have trusted on ML for predicting optimal installation of energy harvesting systems HeliyonReview1 ; PV_forecast_EAAI ; HeliyonReview2 and consequently on predicting energy generation HeliyonPV ; ML_predictUSA ; PV_forecast_EAAI . Similarly, fault detection of these systems were also anticipated thanks to automated models DefectVisual ; DefectEAAI . Others delved into the manufacturing variables of their specific structures ML_structure ; Power_predict ; HeliyonPerov ; ML_OSC . In the same way, there is incipient research on the analysis of stability and durability of these devices PerovBig ; Lifetime ; PEDOT_stability ; PerovML .
@@
In sight of all these developments, this work explores different ML models to encode the PCE performance of polymeric OSCs based on multilayer ITO/PEDOT:PSS/P3HT:PCBM/Al. This permits producing models that reliable learn from the PCE evolution over time. Furthermore, several dependencies on the variables involved in the manufacturing PhysicaB_Dopant_study , but also on the environmental conditions DegradationVariables ; ThermalAnalisis ; DegradationHumidity , are also learnt by these models. To that aim, we have used an hyper-optimizer software, ROBERT ROBERT , that permits benchmarking with multiple ML models and so that to produce extensive and standardized results, which are generated with optimal methods. Our dataset consists of PCE measurements of 45 OSC devices with variability in up to 7 variables, acquired for more than 180 days. This means a database with 996 entries. The accuracy metrics demonstrate the validity of these models to learn and represent the temporal behavior of our OSCs, with determination coefficient up to R2=0.96, and root mean squared error (RMSE), sum of squared error (SSE) and mean absolute error (MAE) bounded around 1% of the target value, that is, the PCE. Additionally, traditional Bayesian regression models sustained by LMS have been introduced in order to confirm the advantages of the ML-automated protocols against these classical approaches. While these conventional methodologies such as PVBayes ; PVStatistical ; PVStatistical2 may yield satisfactory results in this domain, our approach demonstrates that the potential of computational learning modeling offers a much broader and comprehensive solution compared to those based on classical statistical models, showcasing, among other benefits, increased robustness and reliability for:
  • Capturing multivariate relationships between variables, encompassing both linear and non-linear relationships.
@@
  • Learning from an entire dataset, not solely from a specific OSC device as in classical regression.
@@
  • Predicting optimal variables to fabricate the OSC that maximizes the PCE and/or its stability over time.
In contrast ML models, typical statistical algorithms are unable to predict; instead, they utilize estimations to extrapolate target values to immediately subsequent time instances, which must also be exclusive to a particular OSC rather than to an entire database with variability across different variables.
@@
    Assessment of optimal ML models that learn the PCE behavior over time for polymer-based OSC devices with ITO/PEDOT:PSS/P3HT:PCBM/Al structure, considering various fabrication variables and environmental measurement conditions.
@@
    Identification of an optimal ML model capable of accurately predicting the PCE behavior of unseen OSC devices, unknown during the training and validation phases.
@@
    Analysis of dependencies among the dataset variables to establish influence on the performance of the OSCs devices and their fabrication.
@@ 2.1 Organic Solar Cells
The OSCs characterized in this work are manufactured through the spin-coating technique, which involves the deposition of overlapped polymeric thin films. Polymers are deposited in solution, then a process of rotation followed by heating aids in removing the solvent, hence forming the layer. The structure of the devices was as follows: ITO/PEDOT:PSS/P3HT:PCBM/Al, where each layer refers to:
@@
As a substrate, a glass with a 60 nm thick semi-transparent ITO layer was used. The substrates are placed in the spinner with the ITO layer facing upwards so that the thin layers can be deposited on it. The first layer deposited was the PEDOT:PSS film. PEDOT:PSS was deposited at room temperature at a rotation speed of 6000 rpm for 60 seconds. The remaining solvent was removed by heating at 150C for ten minutes. Once the first layer was deposited, the active layer composed by the P3HT:PCBM polymer blend was applied. This was done at 300 rpm for three minutes. It was then dried at 80C for one hour.
The final layer deposited on the devices is the aluminium film. Aluminium is not deposited using the spin-coating technique but rather by metal evaporation in a vacuum chamber. The equipment consists of a high vacuum chamber, two vacuum pumps, one of which is a high vacuum pump, and a power supply to provide the necessary current to evaporate the aluminium. Once the samples are placed inside the chamber, vacuum is achieved, reaching pressures of 10-6 mbar. Once this pressure is reached, aluminium is evaporated by Joule effect.
@@ 2.2 OSC database
As initially mentioned in Section 1, our dataset consists of PCE measurements of 45 OSC devices where up to 7 variables are registered for more than 180 days. Electrical parameters (J-V curves) of these OSCs are acquired along with climate conditions. The acquisition process comprises a Keithley-2400 equipment that acts as source generator for the voltage sweep as well as recorder of the generated PV current. The J-V curves have to be measured under light conditions (100 mW/cm2, AM 1.5G, and 25C) generated by a solar simulator Newport xenon arc lamp and an AM 1.5G filter. The electrical characterization is completed by determining the characteristic electrical parameters, including the short-circuit current density (Jsc), the open-circuit voltage (Voc), the maximum power point (Pmpp), the fill factor (FF), and the PCE. Obtaining these parameters allows to extract the PCE value as follows:
@@
The block diagram of the equipment for the acquisition system is presented by Figure 1. This permitted obtaining a database with 996 entries. The variables and their units are presented by Table 1.
Figure 1: Block diagram of the acquisition system used to generate the OSC database.
Table 1: Variable details of the dataset.
@@
As an example, Figure 2 presents an OSC contained in the database, for which its current density (J) data have been periodically acquired against voltage (V) over 78 days. Specifically, Figure 2 displays the voltage range on the X-axis from 0 to 0.5 V, while Figure 2 adjusts the scale to show the range from 0 to 0.25 V, allowing for a clearer observation of the evolution of the J-V curve as the days pass by. The degradation principles of these organic devices suggest that over time, these curves, from which maximum power and energy conversion efficiency are obtained, should consistently decline. However, it can be observed in Figure 2 that there are certain temporal points between days 20, 26, 47 and 69, where the current density associated with specific voltages remains close to, or even exceeds, the current density associated with those same voltages but at earlier temporal points. This outline confirms that rather than time, there are other variables influencing the degradation of these devices, which implies that a specific OSC could exhibit slightly better performance in the future under certain conditions.
@@
In the same manner, Figure 3 compares the temporal evolution of three different OSCs in terms of their normalized PCE values, ranging from 0 to 1, for more than 180 days. Once again, it is confirmed that time leads to the degradation of the device’s conversion efficiency. However, it also emerges that differences in the fabrication of the various OSCs, as well as the environmental conditions during their measuring, may influence the trend followed by the PCE over time.
@@
2.3 Regression Problem
As already stated, this study aims to automatically model the degradative behavior exhibited by manufactured OSCs using ML approaches. It is clear that we deal with a regression problem, for which recent literature has demonstrated that is reliably and efficiently solved by computational learning. However, in order to establish further comparisons that highlight and reinforce the abilities of AI-based protocols, a preliminary step is carried out to present regression fitting using traditional LMS-supported Bayesian methods.
Regression is a method for estimating the relationship between a response or output variable and one or more predictor or input variables. Linear and non-linear regression serve as estimators of values between observed data points. From that starting point, a regression model relates response data to predictor data through one or more coefficients. Then a fitting algorithm is needed to calculate some model coefficients from a set of input data. In this sense, parametric fitting involves finding those coefficients, assuming that the data are statistical composed by deterministic and random components. Therefore, the parametric model estimates the deterministic component, whereas the random component is typically described as the error. Considering the model as a coefficient function of the independent variable, the error encodes random fluctuations around a Gaussian probability distribution. In many scopes, the goal is to minimize that error, classically addressed by means of LMS fitting approaches, such as: linear least squares, weighted least squares, robust least squares, non-linear least squares, etc.

### v3

# General Machine Learning Models for Interpreting and Predicting Efficiency Degradation in Organic Solar Cells

@@ Abstract
Within the range of renewable energy sources, photovoltaic (PV) plays a crucial role in addressing the exponential growth in global demand during this decade. Organic solar cells (OSCs) emerge as a promising alternative to silicon-based PVs due to their low cost, lightweight, and sustainable production. Despite achieving power conversion efficiencies (PCEs) exceeding values of 20%, OSCs still face challenges in stability and efficiency. In this context, recent advances in manufacturing, artificial intelligence (AI)-driven and machine learning (ML) models offer opportunities to optimize and screen OSCs for greater sustainability and commercial viability, thus potentially reducing costs while ensuring stable and long term performance. This work presents a set of optimal ML models to represent the temporal degradation suffered by the PCE of polymeric OSCs with a multilayer structure ITO/PEDOT:PSS/P3HT:PCBM/Al. To that aim, we generated a database with 166 entries with measurements of 5 OSCs, which includes up to 7 variables regarding both the manufacturing process and environmental conditions for more than 180 days. Then, we relied on a software framework that brings together a conglomeration of automated ML protocols that execute sequentially against our database by simply command-line interface. This easily permits hyper-optimizing the ML models through exhaustive benchmarking so that optimal models are obtained. The accuracy for predicting PCE over time reaches values of the coefficient determination (R2) widely exceeding 0.90, whereas the root mean squared error (RMSE), sum of squared error (SSE), and mean absolute error (MAE) are significantly low. Additionally, we assessed the predictive ability of the models using an unseen OSC as an external set. For comparative purposes, classical Bayesian regression fitting based on non-linear least squares (LS) are also presented, which only perform sufficiently for univariate cases of single OSCs.
@@
Organic solar cells, power efficiency degradation, multilayer structure, Machine Learning, Regression.
journal: Expert Systems with Applications
@@
[label3]organization=Department of Inorganic Chemistry, Institute of Chemical Synthesis and Homogeneous Catalysis (ISQCH), CSIC, University of Zaragoza,addressline=Pedro Cerbuna 12, city=Zaragoza, postcode=50009, country=Spain
@@ 1 Introduction
Electricity consumption has evolved exponentially over the last three decades, increasing from a total estimate of 15277 to 29479 TWh between 2000 and 2023, respectively Rit2023 . One of the immediate consequences has been reflected in the price of this energy. Simultaneously, in the technological realm, there has been a notable shift in electricity generation towards renewable sources. Particularly, the maturity of conventional photovoltaic (PV) cells and the emergence of new organic and hybrid materials have made it possible to get significant improvements in PV generation and its power conversion efficiency (PCE) NREL . However, even though the global consumption of PV energy has increased from 1.08 to 1629.9 TWh between 2000 and 2023 Rit2023 , its generation still falls short of meeting current demands. Hence, there are many challenges and open lines for improving the generation and efficiency of PV cells.
@@
In this context, organic solar cells (OSCs) have emerged as a promising alternative to silicon-based solar cells since, amongst other benefits, their production involves low-temperature manufacturing methods Seo2015 and reduced carbon footprint. Their appeal also lies in the ease of processing, low cost, flexibility, and lightweight Yeh2013 . Moreover, the evolution of their PCEs has been much superior to classical technologies and it has been on par with other considered emerging technologies, presenting efficiencies that exceed values of 20% veintePCE . Even though the commercialization of OSCs still faces various challenges, including stability and efficiency, some cost studies PVCost1 ; PVCost2 ; FER_Heliyon indicate that once these obstacles are overcome, OSCs can be manufactured at a cost lower than one dollar per peak Watt Bra2005 .
@@
The integration of machine learning (ML) before transferring to industrial implementation is being extensively tested these days. For instance, within the field of cheminformatics, it is of paramount importance in the design of new materials. Numerous studies have demonstrated the potential of deep learning algorithms to analyze complex chemical data, thereby accelerating the process of discovering, for example, drugs Dalmau2024 , as well as identifying promising compounds with improved properties Lus2013 ; Sanosa2024 . Some early studies Han2013 ; Rup2012 laid the foundations for the application of ML models in predicting molecular properties, highlighting their great strengths. More recently, the deployment of artificial intelligence (AI) techniques with high-performance prediction methods has enabled the emergence of software platforms capable of examining extensive chemical databases to identify potential materials with significant properties Unt2014 ; Wall2015 . Furthermore, researchers moved forward from general chemical structures to the inclusion of PV features Joh2011 , predicting the electrical behavior of PV devices in terms of parameters such as voltage, current, conversion efficiency, short-circuit current, open-circuit voltage, etc. Eib2021 ; Ser2016 ; Tol2019 ; Lop2016 . Currently, there is a wide array of ML and deep learning-based engines Moo2022 ; Eib2021 ; Mal2021 ; Seo2019 ; Sun2019 designed to optimize different target features of optoelectronic devices Miy2021 ; Mah2022a ; Mah2022b .
Other authors concentrated on AI-supported models to propose long lifecycle materials for sustainable energy generation  Vel2023 ; Ju2018 ; Mah2023 . Some studies have used ML for predicting optimal installation of energy harvesting systems  HeliyonReview1 ; PV_forecast_EAAI ; HeliyonReview2 whereas others have focused on predictors of energy generation HeliyonPV ; ML_predictUSA ; PV_forecast_EAAI . Likewise, fault detection of PV systems has been also anticipated thanks to automated models  DefectVisual ; DefectEAAI . Others delved into the manufacturing variables of the specific PV structures ML_structure ; Power_predict ; HeliyonPerov ; ML_OSC . In the same way, there is incipient research on the analysis of stability and durability of these devices PerovBig ; Lifetime ; PEDOT_stability ; PerovML .
@@
In sight of all these developments, this work explores different ML models to encode the PCE performance of polymeric OSCs based on multilayer ITO/PEDOT:PSS/P3HT:PCBM/Al. To that aim, we have used the ROBERT program ROBERT , which automates data curation, screening of ML models, assessment of predictive ability, and feature analysis (for reproducibility details, see C). Our dataset consists of PCE measurements acquired for 180 days, of 5 OSC devices, resulting in a database with 166 entries. Although these devices are encapsulated, the database contains up to seven descriptors, some related to environmental conditions (temperature, hummidity, dew point and pressure) to inspect possible dependencies on their future performance and some others related to their composition and manufacturing (quantities of: solvent in the HTL layer, i.e., PEDOT:PSS; P3HT, PCBM; and volume ratio of P3HT:PCBM). The accuracy metrics demonstrate the validity of these models to learn and represent the temporal behavior of our OSCs, with determination coefficient up to R2=0.96, and low root mean squared error (RMSE), sum of squared error (SSE) and mean absolute error (MAE). Additionally, traditional Bayesian regression models sustained by non-linear least squares have been introduced to confirm the advantages of the ML-automated protocols against these classical approaches. While these conventional methodologies PVBayes ; PVStatistical ; PVStatistical2 may yield satisfactory results in univariate regression domains, our approach demonstrates that computational learning modeling offers a much broader and comprehensive solution compared to those based on classical statistical models. ML models show increased robustness and reliability for:
  • Generating multivariate models to capture PCE behavior over time while ensuring the interpretability of variable influences.
@@
  • Learning from an entire dataset of multiple OSCs, not solely from a specific OSC device as in classical regression.
@@
  • Detecting optimal variables to fabricate the OSCs with optimal PCE and/or its stability over time.
@@
    Assessment of optimal ML models that learn the PCE behavior over time for polymer-based OSC devices with ITO/PEDOT:PSS/P3HT:PCBM/Al structures.
@@
    Identification of an optimal ML model capable of predicting the PCE behavior of unseen OSC devices.
@@
    Feature analysis to establish the influence of variables on the performance of the OSC devices.
@@ 2.1 Organic Solar Cells
The OSCs characterized in this work are manufactured through spin-coating technique, which involves the deposition of overlapped polymeric thin films. Polymers are deposited in solution, then a process of rotation followed by heating aids in removing the solvent, hence forming the layer. The structure of the devices was as follows: ITO/PEDOT:PSS/P3HT:PCBM/Al, where each layer refers to:
@@
As a substrate, a glass with a 60 nm thick semi-transparent ITO layer was used. The substrates were placed in the spinner with the ITO layer facing upwards so that the thin layers could be deposited on it. The first layer was the PEDOT:PSS film. It was deposited at room temperature at a rotation speed of 6000 rpm for 60 seconds. The remaining solvent was removed by heating at 150 C for 10 minutes. Once the first layer was deposited, the active layer composed by the P3HT:PCBM polymer blend was applied. This was done at 300 rpm for 3 minutes. It was then dried at 80 C for 1 hour.
The final layer deposited on the devices was an aluminium film. Aluminium was not deposited using the spin-coating technique but rather by metal evaporation in a vacuum chamber. The equipment consists of a high vacuum chamber, two vacuum pumps, one of which is a high vacuum pump, and a power supply to provide the necessary current to evaporate the aluminium. Once the samples were placed inside the chamber, vacuum was achieved, reaching pressures of 10-6 mbar. When this pressure was reached, the aluminium evaporated by Joule effect.
@@ 2.2 OSC database
As initially mentioned in Section 1, our dataset consists of PCE measurements of 5 OSC devices, where up to 7 variables were registered for more than 180 days. Electrical parameters (J-V curves) of these OSCs were acquired along with climate conditions. The acquisition process comprises a Keithley-2400 equipment that acts as source generator for the voltage sweep as well as recorder of the generated PV current. The J-V curves were measured under light conditions (100 mW/cm2, AM 1.5 G, and 25 C) generated by a solar simulator Newport xenon arc lamp and an AM 1.5G filter. The electrical characterization was completed by determining the characteristic electrical parameters, including the short-circuit current density (Jsc), the open-circuit voltage (Voc), the maximum power point (Pmpp), the fill factor (FF), and the PCE. Obtaining these parameters allowed to extract the PCE value as follows:
@@
The block diagram of the equipment for the acquisition system is presented in Figure 1. This permitted obtaining a database with 166 entries. The variables and their units are presented in Table 1. Notice that, in order to get further insights on the future performance of the OSCs, environmental conditions at laboratory have also been measured, despite the devices were encapsulated.
Figure 1: Block diagram of the acquisition system used to acquire the OSC database.
Table 1: Detail of the dataset: Manufacturing variables and environmental conditions at laboratory.
@@
As an example, Figure 2 presents an OSC device contained in the database, for which its current density (J) data have been periodically acquired against voltage (V) over time. Specifically, Figure 2 displays the voltage range on the X-axis from 0 to 0.5 V, while Figure 2 adjusts the scale to show the range from 0 to 0.25 V, allowing for a clearer observation of the evolution of the J-V curve when days go by (different colors of the legend). The degradation principles of these organic devices suggest that over time, these curves, from which maximum power and energy conversion efficiency are obtained, should consistently decay.
@@
Likewise, Figure 3 compares the temporal evolution of three different OSCs in terms of their normalized PCE values, ranging from 0 to 1, for more than 180 days. Please note that normalization is used for visualization purposes, as typically used in this field Wurfel when analyzing PCE decay. Once again, it is confirmed that time leads to the degradation of the device’s conversion efficiency. However, it also emerges that differences in the fabrication of the OSCs, as well as the environmental conditions during their measuring, may influence the trend followed by the PCE over time.
@@
2.3 Regression Problem with Classical Approaches
This study aims to automatically model the degradative behavior exhibited by manufactured OSCs using ML approaches, which is a regression problem. In order to establish further comparisons that highlight and reinforce the abilities of ML-based protocols, a preliminary step was carried out to present regression fitting using traditional LS-supported Bayesian methods.
Regression is a method for estimating the relationship between a response or output variable and one or more predictor or input variables. Linear and non-linear regression serve as estimators of values between observed data points. From that starting point, a regression model relates response data to predictor data through one or more coefficients. Then a parametric fitting algorithm is needed to calculate some model coefficients from a set of input data. Therefore, the parametric algorithm estimates deterministic components, whereas random components are typically described as the error. Considering the model as a coefficient function of the independent variable, the error encodes random fluctuations around a Gaussian probability distribution. In many scopes, the goal is to minimize that error, classically addressed by means of least squares (LS) fitting approaches, such as: linear least mean squares, weighted least squares, robust least squares, non-linear least squares, etc.